OpenStudy (anonymous):

Solve p(x)=x4 + 81x2 in comma seperated list

OpenStudy (anonymous):

p(x)=x^4 + 81x^2 p(x)=t^2 + 81t, t = x^2 solve using quadratic rules.

OpenStudy (anonymous):

\[P(x)=x ^{4 }+81x ^{2}\]

OpenStudy (anonymous):

exactly.

OpenStudy (anonymous):

is that what they consider factor p completely...man I should have paid more attention

OpenStudy (anonymous):

p(x)=x^4 + 81x^2 p(x)=t^2 + 81t, t = x^2 p(x)=t(t+81) t=0, t=-81 x^2=0, x^2=-81 x^2=0, no other solutions.

OpenStudy (anonymous):

x=0 is the answer to p(x) = 0

OpenStudy (anonymous):

did i do this one right then...\[p(x)=x ^{4}+6x ^{2}+9\]

OpenStudy (anonymous):

i got \[(x ^{2}+3)(x ^{2}+3)\]